# Ensemble Kalman Filtering Meets Gaussian Process SSM for Non-Mean-Field and Online Inference

### v3

@@ VI-D Online Learning and Inference Using OEnVI
We evaluate the three online learning methods using ${}^{\circledR}$ data, a dataset previously utilized in [55], showcasing dynamics akin to recurrent switching linear dynamical systems [57]. The simulated trajectory faithfully reproduces the layout of a ${}^{\circledR}$ track, as depicted in Fig. 5a. We train these three methods with $2000$ observations and test them with $500$ observations, both simulated from ${\mathbf{y}}_{t}=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t}$ where $\bm{C}$ is a 10-by-2 matrix generated randomly. The SVMC and VJF algorithms are implemented using the code publicly provided online [^5] https://github.com/catniplab/svmc [^6] https://github.com/catniplab/vjf.
Figs. 5a5c depict the true states (in blue) and the latent states (in red) inferred by the three methods. The results clearly indicate that EnVI and SVMC swiftly captured the real state and maintained accuracy, while VJF faced challenges, primarily due to its difficulty in optimizing the parameters in the inference network. The detailed comparison with SVMC, illustrated in Figs. 5d5f, highlights the superior accuracy of OEnVI in both inference and prediction of latent states. These empirical findings emphasize the efficacy of OEnVI, particularly in its use of EnKF to approximate the variational distribution, demonstrating its advancement over VJF using inference networks, and SVMC using PF methods.
@@ VII Conclusion
In this paper, we have introduced EnVI, a novel NMF algorithm tailored for GPSSMs, which integrates EnKF into a variational inference framework. Additionally, we have presented an extended online version, OEnVI. Both algorithms eliminate the necessity for heavy parameterization like inference networks and shape the variational distribution over latent states through the model-based EnKF. Leveraging the inherent differentiable nature along with the modern automatic differentiation tools, the proposed EnVI and OEnVI can enhance efficiency and algorithmic robustness while improving learning and inference performance compared to existing methods. Detailed analysis and fresh insights for the proposed algorithms are provided to enhance their interpretability. Empirical experiments conducted on diverse real and synthetic datasets, evaluating filtering, prediction, and dynamics learning performance, unequivocally support the effectiveness of the proposed methods.
@@ Appendix A Model Evidence Lower Bound Evaluations
Under the approximations that $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, $\forall t$, computing the difference between term 1 and term 4 in the ELBO (Eq. (21)) yields an approximation to an expected log-likelihood, i.e.,
$\operatorname{term~{}1}-\operatorname{term~{}4}\approx\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{1:T}|{\vec{\mathbf{u}}})\right]$ (37)
@@
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]$ (38a)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}\right]$ (38b)
$\!\approx\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}\right]$ (38c)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]$ (38d)
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right],$ (38e)
where Eq. (38c) holds because of the approximation of $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ and Eq. (38e) is derived straightforwardly by applying Bayes’ theorem. ∎
@@
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (39) can be analytically evaluated, see Eq. (27), due to the Gaussian prediction distribution, see Eq. (23), and the linear emission model.
@@ Appendix B Proof of Proposition 2
Under the approximation that $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, we have the log-likelihood term
@@ D-A Evidence Lower Bound (ELBO)
$\!\!\!\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})=\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}$ (43)
$\quad=\int\int q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\mathrm{d}{\vec{\mathbf{x}}}\mathrm{d}{\vec{\mathbf{f}}}$
$\quad=\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\text{ELBO: }\mathcal{L}}+\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}\right]}_{\text{KL divergence}}$
@@ E-A OEnVI: Online Learning Results on Linear Gaussian SSMs
  • The baseline in Table III is the RMSE between the noisy observations ${\mathbf{y}}_{1:T}$ and the latent states ${\mathbf{x}}_{1:T}$
TABLE III: State inference performance (RMSE) of EnVI and OEnVI
@@
Figure 6: OEnVI on learning and inference in linear Gaussian SSMs
@@
Figure 7: OEnVI on learning and inference in linear Gaussian SSMs
@@ E-D Online Learning and Inference with OEnVI
Figure 10: Kink transition function learning using Online EnVI

### v5

@@ VI-D Online Learning and Inference Using OEnVI
We evaluate the three online learning methods using NASCAR® data, a dataset previously utilized in [59], showcasing dynamics akin to recurrent switching linear dynamical systems [61]. The latent state trajectory faithfully reproduces the layout of a NASCAR® track, as depicted in Fig. 5a. We train these three methods with $2000$ observations and test them with $500$ observations, both generated from ${\mathbf{y}}_{t}=\bm{C}{\mathbf{x}}_{t}+\mathbf{e}_{t}$, where $\bm{C}$ is a 10-by-2 matrix generated randomly (strictly following the settings of SVMC [59]), and $\mathbf{e}_{t}\!\sim\!{\cal N}(\bm{0},0.1^{2}\bm{I}_{2\times 2})$. The SVMC and VJF algorithms are implemented using the code publicly provided online [^5] https://github.com/catniplab/svmc [^6] https://github.com/catniplab/vjf.
Figs. 5a5c depict the true states (in blue) and the latent states (in red) inferred by the three methods. The results clearly indicate that EnVI and SVMC swiftly captured the real state and maintained accuracy, while VJF faced challenges, primarily due to its difficulty in optimizing the parameters in the inference network. The detailed comparison with SVMC and VJF, illustrated in Figs. 5d5f, highlights the superior accuracy of OEnVI in both inference and prediction of latent states. These empirical findings emphasize the efficacy of OEnVI, particularly in its use of EnKF to approximate the variational distribution, demonstrating its advancement over VJF using inference networks, and SVMC using PF methods.
@@ VII Conclusion
In this paper, we have introduced EnVI, a novel NMF algorithm tailored for GPSSMs, which integrates EnKF into a variational inference framework. Additionally, we have presented an extended online version, OEnVI. Both algorithms eliminate the necessity for heavy parameterization like inference networks and shape the variational distribution over latent states through the model-based EnKF. Leveraging the inherent differentiable nature along with the modern automatic differentiation tools, the proposed EnVI and OEnVI can enhance efficiency and algorithmic robustness while improving learning and inference performance compared to existing methods. Detailed analysis and fresh insights for the proposed algorithms are provided to enhance their interpretability. Empirical experiments conducted on diverse real and synthetic datasets, evaluating filtering, prediction, and dynamics learning performance, unequivocally support the effectiveness of the proposed methods. Future research will explore efficient learning and inference techniques for complex time-varying dynamical systems.
@@ Appendix A Model Evidence Lower Bound Evaluations
Under the approximations that:
  • 1)
    $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$,
  • 2)
    $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$,
computing the difference between term 1 and term 4 in the ELBO (Eq. (37)) yields the expected log-likelihood, i.e.,
$\operatorname{term~{}1}-\operatorname{term~{}4}=\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\log p({\mathbf{y}}_{1:T}|{\vec{\mathbf{u}}})\right]$ (38)
@@
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})}\right]$
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}}{q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}}\right]$
$\!\approx\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\!\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\underbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{assumption 2)}}}\underbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{assumption 1)}}}}\right]$
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}},{\vec{\mathbf{x}}})}\left[\sum_{t=1}^{T}\log\frac{p({\mathbf{y}}_{t},{\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})}{p({\mathbf{x}}_{t},{\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1},{\mathbf{y}}_{t})}\right]$
$\!=\!\mathbb{E}_{q({\vec{\mathbf{u}}})}\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right],$ (39)
where the last line of Eq. (39) is derived straightforwardly by applying Bayes’ theorem. ∎
@@
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (40) can be analytically evaluated using EnKF, see Eq. (27), due to the Gaussian prediction distribution, see Eq. (25), and the linear emission model.
@@ Appendix B Proof of Proposition 2
With the filtering distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$, we have the log-likelihood term
@@ D-A Evidence Lower Bound (ELBO)
$\log p({\vec{\mathbf{y}}}|{{\bm{\theta}}})=\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}$ $=\int\int q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\mathrm{d}{\vec{\mathbf{x}}}\mathrm{d}{\vec{\mathbf{f}}}$ (44)
$=\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{p({\vec{\mathbf{y}}},{\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\right]}_{\text{ELBO: }\mathcal{L}}+\underbrace{\mathbb{E}_{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}\!\left[\log\frac{q({\vec{\mathbf{x}}},{\vec{\mathbf{f}}})}{p({\vec{\mathbf{x}}},{\vec{\mathbf{f}}}|{\vec{\mathbf{y}}})}\right]}_{\text{KL divergence}}$
@@ E-A OEnVI: Online Learning Results on Linear Gaussian SSMs
  • The baseline in Table III and Table IV is the RMSE between the noisy observations ${\mathbf{y}}_{1:T}$ and the latent states ${\mathbf{x}}_{1:T}$.
  • Table III, and Figs. 7 and 8 present the results obtained using our proposed methods, EnVI and OEnVI, applied to learning the linear Gaussian SSM with the identity coefficient matrix $\bm{C}=\bm{I}_{4\times 4}$.
  • The results in Table IV and Fig. 6 show that our proposed methods, EnVI and OEnVI, can still accurately filter the true latent states. However, it is noteworthy that the matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$ should be “full-column rank” if the main task is to perform filtering; otherwise, there is a high probability that some latent dimensions cannot be inferred accurately. Intuitively speaking, if we aim to recover latent state ${\mathbf{x}}_{t}$, it is only possible when the corresponding observation ${\mathbf{y}}_{t}$ contains “sufficient information” about ${\mathbf{x}}_{t}$. For the task of predicting the sequence ${\mathbf{y}}_{t}$ for $t=T+1,T+2,\ldots$, it is typically not necessary to focus as much on the physical meaning and accuracy of the latent states.
TABLE III: State inference performance (RMSE) of EnVI and OEnVI (with emission coefficient matrix $\bm{C}=\bm{I}_{4\times 4}$)
@@
TABLE IV: State inference performance (RMSE) of EnVI and OEnVI (with emission coefficient matrix $\bm{C}\in\mathbb{R}^{4\times 4}$ generated randomly )
KF 0.7799 1.0302 1.1086 1.3682 1.4355 1.1277 0.7517 1.1239 1.0911 1.1062
EnVI 1.8413 / / / / / / / / 2.3356
OEnVI 1.4508 1.5902 1.3763 3.0673 1.2875 2.1743 1.2470 3.6032 2.8966 2.1546
Baseline 24.10 100.49 211.82 272.88 303.85 422.38 547.60 741.17 847.08 428.24
Figure 6: EnVI (top) & OEnVI (bottom) on state inference in linear Gaussian SSM, with the emission coefficient matrix $\bm{C}\in\mathbb{R}^{4\times 4}$ generated randomly. The RMSE of the latent state estimates for KF, EnVI, and OEnVI are 0.7799, 1.8413, and 1.4508, respectively; the RMSE between the observations and the latent states is 24.10.
@@
Figure 7: OEnVI on learning and inference in linear Gaussian SSMs with $\bm{C}=\bm{I}_{4\times 4}$
@@
Figure 8: OEnVI on learning and inference in linear Gaussian SSMs with $\bm{C}=\bm{I}_{4\times 4}$
@@ E-D Online Learning and Inference with OEnVI
Figure 11: Kink transition function learning using OEnVI